OpenStudy (anonymous):

If y is a negative integer, which of the following has the least value? cube root of y^2 100^y π/y 1/y^2 1/y^3

OpenStudy (lgbasallote):

it would be a lot easier if you let y be a number..let's say -1 so y = -1 \[\large \sqrt[3]{y^2} \implies \sqrt[3]{(-1)^2} \implies \sqrt[3]{1} \implies 1\] \[\large 100^y \implies 100^{-1} \implies \frac{1}{100}\] \[\large \frac{\pi}{y} \implies \frac{\pi}{-1} \implies -\pi\] \[\large \frac{1}{y^2} \implies \frac{1}{(-1)^2} \implies \frac 11 \implies 1\] \[\large \frac{1}{y^3 }\implies \frac{1}{(-1)^3} \implies \frac{1}{-1} \implies -1\] so which do you think is the smallest one?

OpenStudy (anonymous):

the 2nd one